OpenStudy (anonymous):

Use the Comparison Theorem (CT) to determine whether this integral is convergent or divergent [latex] \[\int\limits_{1}^{\infty}\frac{33(2+e^{-x})}{x}dx\]

OpenStudy (anonymous):

Ahh!Messed it up. One sec.

OpenStudy (zepp):

@FoolForMath

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty}\frac{33(2+e^{-x})}{x}dx\]

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

I said this is divergent, and the comparison I made was with this: \[\int\limits\limits_{1}^{\infty}\frac{33(2+e^{-x})}{1/x}dx\]

OpenStudy (anonymous):

Because: \[x \ge\frac{1}{x}\] on [1,infinity]

OpenStudy (anonymous):

I want to clarify with you guys to see if it is a right or wrong comparison.

OpenStudy (anonymous):

So basically, what I did was I solved for the one with "1/x"

OpenStudy (turingtest):

I don't think that's a good comparison, but I'm struggling to see why we even need the comparison test...

OpenStudy (anonymous):

Because it asks to use the CT.

OpenStudy (anonymous):

Why is it not a good comparison? Is not the denominator term (1/x) smaller than the term (x)

OpenStudy (turingtest):

right, so the integrand is larger for 1/x than with x int the denom so proving that it diverges with 1/x does not prove that it diverges with x we need to find a smaller integrand that also diverges

OpenStudy (anonymous):

I know: \[\frac{e^-x}{x}\]

OpenStudy (turingtest):

\[\frac{33(2+e^{-x})}{x}\le\frac{33(2+e^{-x})}{\frac1x}\]for \(x\ge1\)

OpenStudy (anonymous):

And this has to be smaller.

OpenStudy (anonymous):

Would my suggestion work above?

OpenStudy (anonymous):

\frac{e^-x}{x}

OpenStudy (turingtest):

but\[\frac{e^{-x}}x\le\frac{e^{-x}}{\frac1x}=xe^{-x}\]so by putting 1/x in the denom we are making it bigger

OpenStudy (anonymous):

No I meant that see if the following integral is divergent: \[\int\limits_{1}^{\infty}\frac{e^-x}{x}\]

OpenStudy (anonymous):

Because surely, this is smaller than the integral in question.

OpenStudy (turingtest):

but that integral is not divergent

OpenStudy (anonymous):

Really?

OpenStudy (turingtest):

not sure if I can prove that... I'd have to think about it....

OpenStudy (turingtest):

but your integral is so obviously divergent for other reasons, I just think looking to use the comparison test is overkill

OpenStudy (anonymous):

Unfortunately, I need a comparison. I just realized that the one I came up with will be convergent. Ahhhh....

OpenStudy (turingtest):

oh now I see how to prove the convergence of\[\int_1^{\infty}\frac{e^{-x}}{x}dx\]not that that helps you...

OpenStudy (turingtest):

Let me at least make a quick case for divergence of your integral my way; maybe that will help...

OpenStudy (anonymous):

How about: \[\frac{e-^{x}}{x^2}\]

OpenStudy (turingtest):

also convergent

OpenStudy (turingtest):

if \[\frac{e^{-x}}x\]converges on [1,infty] then so does \[\frac{e^{-x}}{x^2}\]because the second is less than the first for all x in that interval you can tell that just by looking

OpenStudy (anonymous):

It works!

OpenStudy (anonymous):

I think it works...

OpenStudy (turingtest):

My view: in your integral\[\int\limits_{1}^{\infty}\frac{33(2+e^{-x})}{x}dx=33[2\int\limits_{1}^{\infty}\frac1xdx+\int\limits_{1}^{\infty}\frac{e^{-x}}xdx]\]the first integral clearly diverges, though the second does not how you want to apply the comparison test I don't know...

OpenStudy (anonymous):

OHH! Ok I think I will use that strategy to explain why I do not use comparison test.

OpenStudy (anonymous):

What you said above I meant

OpenStudy (turingtest):

there may be a way to use it, but if so it seems trivial.... good luck to any who see how though, back to work for me :) Good luck!

OpenStudy (anonymous):

Thanks

OpenStudy (turingtest):

welcome